OpenStudy (anonymous):

2+4i What is the product?

OpenStudy (anonymous):

i believe you are missing something

OpenStudy (anonymous):

product? where are you multipliying?

OpenStudy (anonymous):

Product=2 numbers multiplied. We have, 2+4i only 1 complex number. PARAMETERS MISSING.

OpenStudy (anonymous):

BBB83 - are you trying to multiply 2 and 4i

OpenStudy (anonymous):

Lol o.o but it's given 2+4i so I am not sure how that is correct.

OpenStudy (anonymous):

(2+4i)(2-4i) = 4 - 16i^2 = 4 + 16 = 20

OpenStudy (anonymous):

sorry i forgot to add the complex conjugate 2-4i

OpenStudy (anonymous):

That's how the complex number work, always go in pair!

OpenStudy (anonymous):

(2+4i)(2-4i) Which is same as, (a+b)(a-b) So, \[a ^{2}-b ^{2}\] Which is, \[2^{2}-(4i)^{2}\] 4-(-16) 4+16=20